A certain green light has a frequency of 6.0 x 1014 Hz.
What is the energy content of one quantum of the light?
1
Expert's answer
2020-11-09T13:59:50-0500
"E = h\nu = h {c\over \lambda}""E = 6.626*10^{-34} {3*10^8\over 6*10^{14}} = 3.313*10^{-40} J"